# Shifting AI Efficiency From Model-Centric to Data-Centric Compression

### v1

@@ 3.2 Training and Inference Targets - How Token Compression Benefits?
$\bullet$ Data augmentation techniques have been widely adopted to enrich training datasets by introducing variability that enhances robustness and informativeness (Autoaugment, ). In computer vision, mixing or combining image tokens creates novel representations that elevate training effectiveness (mixup, ; cutmix, ). This strategy has also been extended to synthetic datasets, where adaptive augmentation controls the informativeness of generated images (DCC, ; DSA, ; NCFM, ). Analogously, in natural language processing, augmenting text tokens through synonym replacement (wei2019eda, ), contraction expansion (coulombe2018text, ), back-translation (chen2020mixtext, ), and reformulation (hao2025maga, ), supporting better generalization.
$\bullet$ Token selection focuses on filtering out low-quality tokens to refine training data quality (rho1, ; lee2022deduplicating, ; penedo2023refinedweb, ; wenzek2019ccnet, ; gao2020pile, ; li2024datacomp, ). Common approaches include rule-based heuristics (raffel2020exploring, ; penedo2024fineweb, ), deduplication methods (lee2022deduplicating, ; penedo2023refinedweb, ; abbas2023semdedup, ), and scoring strategies leveraging large language models (wenzek2019ccnet, ; gao2020pile, ; li2024datacomp, ; wettig2024qurating, ; sachdeva2024train, ).
@@
(II) Increasing Training Efficiency Token compression directly reduces the token length processed during training, addressing critical challenges associated with scaling large models (Bolya2023:ToMe, ; Choudhury2024:RLT, ; Shang2024:LLaVA-PruMerge, ; xing2024pyramiddrop, ). For Transformer architectures with sequence length reduced from $n$ to $m$ ($m<n$), the computational and memory benefits can be quantified as:
@@
(I) Decreasing Computational Complexity Following patterns established in training, token compression achieves quadratic speedup in inference computations. Notably, many non-parametric compression methods (Bolya2023:ToMe, ; Ju2024:Turbo, ; Chen2024:FastV, ) can be directly integrated into inference without additional training or architectural modifications, enabling immediate benefits across various domains (zhang2024sparsevlm, ; wen2025stop, ).
(II) Reducing Memory Usage Token compression optimizes memory efficiency through two mechanisms: (i) computing memory reduction following the linear scaling pattern shown in training, and (ii) KV cache optimization for large language models (li2024snapkv, ; cai2024pyramidkv, ; wan2024look-m, ; wan2025meda, ). During autoregressive generation, each layer caches key and value states for attention computation, with memory growing with sequence length. For a sequence of length $n$ compressed to length $m$, with $L$ layers and hidden dimension $d$, the KV cache memory reduction is:
@@
These benefits are particularly crucial for real-time interactive systems, including UI agents (tang2025guisurvey, ), autonomous driving (gao2021autonomous, ), and embodied AI (duan2022embodied, ), where efficient processing of continuous inputs under resource constraints is essential.
@@ 3.3 Compelling Advantages - Why Token Compression Matters?
  • Quadratic Gains: The $\mathcal{O}(n^{2})$ computation complexity of widely used self-attention indicates token compression can bring significant benefits in computation..
@@ 4.1 Performance Degradation
Methodological Bottlenecks. Attention scores play a crucial role across existing token compression approaches. For example, [CLS] token attention scores are used to select key visual tokens (qiang2022attcat, ; haurum2023tokens, ; zhang2024cls, ; wang2024cls, ; han2024rethinking, ; yang2024visionzip, ; liu2025compression, ), while cross-modal guidance (jiang2022trips, ; cao2023pumer, ; shi2023crossget, ; chen2024image, ; xing2024pyramiddrop, ; zhang2024sparsevlm, ) relies on text-vision attention scores. But are attention scores truly reliable for deciding which tokens to keep? Recent work (zhang2024cls, ; wen2025token, ) reveals that attention scores can suffer from position bias. For instance, when using text-vision scores to retain visual tokens, those near the sequence end often get higher weights. In 2D image space, this biases retention toward the lower half or bottom-right corner. Clearly, it’s unrealistic to assume the lower half of all images is more important. Such bias can significantly hurt compression performance. As shown in Figure 2, a comprehensive review and comparison of key studies (jiang2025kind, ; wen2025stop, ; wen2025token, ; Zhang2024:SparseVLM, ; liu2024multi, ; liu2025vidcom2, ) confirms our hypothesis: even well-crafted attention-based methods can underperform simple random pruning or pooling. Detailed analysis is in Appendix B.
Inherent Limitations of Token Compression. Beyond performance degradation from methodological design, does token compression face inherent limitations? Is it universally applicable across various tasks? For multi-modal large language models, (wen2025token, ) shows most existing compression methods underperform on visual grounding tasks, with significant drops on benchmarks like RefCOCO (yu2016modeling, ). In OCR-related parsing (yang2024cc, ; ouyang2024omnidocbench, ), documents with dense layouts yield highly information-rich visual tokens. Compressing these risks severe information loss and degraded performance. Beyond vision, current methods also face inherent limits in other modalities. In Automatic Speech Recognition (ASR) and Automatic Speech Translation (AST) (ardila2019common, ; conneau2023fleurs, ), audio is encoded and then decoded into text using an MLLM (abouelenin2025phi, ; chu2024qwen2, ). Audio tokens in ASR and AST are dense and temporally continuous. Pruning or merging them disrupts this continuity, leading to fragmented recognition or translation. Similarly, translation across languages in the text modality may suffer significant degradation under high compression ratios.
@@ 4.3 Fair Comparison
Rethinking FLOPs and Compression Ratios as Efficiency Metrics. Many token compression methods report speedup by estimating FLOPs reductions or directly using token compression ratios. But do FLOPs or compression ratios truly reflect real acceleration? Our analysis shows that, even with similar compression ratios or FLOPs, methods often vary significantly in runtime latency. Investigating further, we find: (i) Importance-based compression often uses attention scores (Chen2024:FastV, ; zhang2024sparsevlm, ), but this can limit compatibility with efficient attention mechanisms like Flash Attention (dao2022flashattention, ; dao2023flashattention2, ), potentially contributing to the substantial discrepancy between FLOPs and actual runtime latency observed in some approaches. (ii) Some methods pursue high compression via progressive compression across layers (xing2024pyramiddrop, ), adding overhead that offsets the gains from token reduction. Thus, we argue that runtime latency should be prioritized in evaluations, as FLOP or token count reductions don’t always yield real-world speedup.
@@
Token Compression Evaluation: The Benchmarking Gap. Current token compression methods are primarily evaluated using general-purpose benchmarks, which are not tailored to capture the unique challenges of token compression. As a result, some benchmarks—such as ScienceQA (lu2022learn, ) and VizWiz (bigham2010vizwiz, )—paradoxically show improved performance under certain compression settings, or minimal degradation across varying compression ratios. These observations defy intuition and suggest that existing benchmarks may fail to meaningfully reflect the trade-offs introduced by token compression. This discrepancy raises concerns about the validity of current evaluation practices. In particular, benchmarks that fail to penalize information loss introduced by compression may obscure meaningful differences between methods. Additionally, the lack of task diversity and compression-sensitive metrics further limits our understanding of how these methods behave in realistic scenarios. Without dedicated benchmarks, it remains unclear whether observed gains are due to genuine improvements in compression quality or artifacts of misaligned evaluation settings.
@@ 5.2 Dedicated Benchmarks for Token Compression
Given the current limitations in evaluating data-centric token compression methods using general-purpose benchmarks, we envision the development of a dedicated benchmark specifically designed to evaluate them. Such a benchmark should comprehensively span diverse domains—including natural language processing, computer vision, and multi-modal tasks—and incorporate task-specific challenges particularly relevant to token compression, such as optical character recognition (OCR) parsing (ouyang2024omnidocbench, ; zhang2024ocr, ) and automatic speech recognition (ASR) (conneau2023fleurs, ; park2024long, ). Furthermore, it is essential that this benchmark jointly considers both task performance and latency, as both are critical for real-world deployment. A well-rounded benchmark of this nature would enable more rigorous, fair, and holistic evaluation of token compression techniques, ultimately driving progress in this area.
@@ 6.1 Model-Centric Compression as a Superior Alternative
Model-centric compression methods, such as pruning (Network_Pruning, ; JMLR:v12:huang11b, ; Deep_Compression, ; Rao2021:DynamicViT, ; jang2017Gumbel-Softmax, ; Pan2021:IA-RED2, ), quantization (rokh2023comprehensive, ; Wang:HAQ, ; zhou2018adaptive, ; yang2019quantization, ; Lin:AWQ, ; Frantar:GPTQ, ), and knowledge distillation (Hinton:KD, ; zhang2019self_distillation, ; zhang2021self_distillation, ; park2019relational_distillation, ), have long been established as effective techniques for reducing model size and computational cost. Proponents argue that this paradigm is reliable for deployment in resource-constrained environments and maintains performance consistency. For example, pruning techniques such as DynamicViT (Rao2021:DynamicViT, ) dynamically remove uninformative tokens during inference, reducing the computational load by up to 30–40% with minimal impact on accuracy. Proponents of this view claim that this approach achieves substantial speedups without discarding any original data. In contrast, data-centric methods that prune input tokens risk removing critical contextual information, which may degrade performance.
Counterargument. Although model-centric compression is effective, it faces scalability issues as models and datasets grow, requiring costly full retraining and processing of entire inputs. In contrast, data-centric compression reduces input complexity upfront, easing computational burdens. Some data-centric methods update only a small parameter subset (Zhao2024:DyT, ; liu2024Sparse-Tuning, ; Lee2024:DETM, ), while others enable training-free deployment (bolya2023tome, ; Chen2024:FastV, ; li2024snapkv, ). Combining both approaches can improve efficiency without sacrificing accuracy (azeemi2023data, ; kousar2025pruning, ), making data-centric methods a complement to model-centric techniques.
@@ 6.2 Advanced Model Architectures as a more Promising Direction
Another argument against data-centric compression is the continued advancement of model architectures that can inherently handle large datasets and long sequences more efficiently (gu2021efficiently, ; goldstein2024goldfinch, ; peng2023rwkv, ). The development of transformer-based architectures, such as Vision Transformers (ViT, ), Swin Transformers (liu2021swin, ), and large language models like GPT-3 (brown2020language, ), has shown significant improvements in both accuracy and scalability. These architectures integrate advanced techniques, such as hierarchical processing, self-attention mechanisms, and dynamic sparsity, enabling them to process large amounts of data efficiently. For example, Swin Transformers (liu2021swin, ) utilize a window-based self-attention mechanism, which reduces the computational complexity of the standard attention mechanism, making it feasible to scale models to much larger datasets and sequences. Proponents of this view argue that as these advanced models continue to evolve, there may be less need for aggressive input compression, as these models are inherently better equipped to handle large-scale data directly.
Counterargument. Advanced model architectures offer strong performance but demand substantial computational resources, especially during training (llama4, ; yang2025Qwen3, ; OpenAI:GPT-4, ). Data-centric compression reduces computational load early by simplifying input data, enabling more efficient training and inference without sacrificing accuracy. Techniques like token pruning and augmentation preserve or improve performance by focusing on informative data. Combined with advanced architectures, data-centric methods enhance efficiency and maintain high performance (gao2020pile, ; wettig2024qurating, ), making them complementary rather than competitive.

### v2

@@ 3.2 Training and Inference Targets - How Token Compression Benefits?
$\bullet$ Data augmentation techniques have been widely adopted to enrich training datasets by introducing variability that enhances robustness and informativeness Autoaugment . In computer vision, mixing or combining image tokens creates novel representations that elevate training effectiveness mixup ; cutmix . This strategy has also been extended to synthetic datasets, where adaptive augmentation controls the informativeness of generated images DCC ; DSA ; NCFM . Analogously, in natural language processing, augmenting text tokens through synonym replacement wei2019eda , contraction expansion coulombe2018text , back-translation chen2020mixtext , and reformulation hao2025maga , supporting better generalization.
$\bullet$ Token selection focuses on filtering out low-quality tokens to refine training data quality rho1 ; lee2022deduplicating ; penedo2023refinedweb ; wenzek2019ccnet ; gao2020pile ; li2024datacomp . Common approaches include rule-based heuristics raffel2020exploring ; penedo2024fineweb , deduplication methods lee2022deduplicating ; penedo2023refinedweb ; abbas2023semdedup , and scoring strategies leveraging large language models wenzek2019ccnet ; gao2020pile ; li2024datacomp ; wettig2024qurating ; sachdeva2024train .
@@
(II) Increasing Training Efficiency Token compression directly reduces the token length processed during training, addressing critical challenges associated with scaling large models Bolya2023:ToMe ; Choudhury2024:RLT ; Shang2024:LLaVA-PruMerge ; xing2024pyramiddrop . For Transformer architectures with sequence length reduced from $n$ to $m$ ($m<n$), the computational and memory benefits can be quantified as:
@@
(I) Decreasing Computational Complexity Following patterns established in training, token compression achieves quadratic speedup in inference computations. Notably, many non-parametric compression methods Bolya2023:ToMe ; Ju2024:Turbo ; Chen2024:FastV can be directly integrated into inference without additional training or architectural modifications, enabling immediate benefits across various domains zhang2024sparsevlm ; wen2025stop .
(II) Reducing Memory Usage Token compression optimizes memory efficiency through two mechanisms: (i) computing memory reduction following the linear scaling pattern shown in training, and (ii) KV cache optimization for large language models li2024snapkv ; cai2024pyramidkv ; wan2024look-m ; wan2025meda . During autoregressive generation, each layer caches key and value states for attention computation, with memory growing with sequence length. For a sequence of length $n$ compressed to length $m$, with $L$ layers and hidden dimension $d$, the KV cache memory reduction is:
@@
These benefits are particularly crucial for real-time interactive systems, including UI agents tang2025guisurvey ; tang2025gui , autonomous driving gao2021autonomous , and embodied AI duan2022embodied ; yang2025efficientvla , where efficient processing of continuous inputs under resource constraints is essential.
@@ 3.3 Compelling Advantages - Why Token Compression Matters?
  • Quadratic Gains: The $\mathcal{O}(n^{2})$ computation complexity of widely used self-attention indicates token compression can bring significant benefits in computation.
@@ 4.1 Performance Degradation
Methodological Bottlenecks. Attention scores play a crucial role across existing token compression approaches. For example, [CLS] token attention scores are used to select key visual tokens qiang2022attcat ; haurum2023tokens ; zhang2024cls ; wang2024cls ; han2024rethinking ; yang2024visionzip ; liu2025compression , while cross-modal guidance jiang2022trips ; cao2023pumer ; shi2023crossget ; chen2024image ; xing2024pyramiddrop ; zhang2024sparsevlm relies on text-vision attention scores. But are attention scores truly reliable for deciding which tokens to keep? Recent work zhang2024cls ; wen2025token reveals that attention scores can suffer from position bias. For instance, when using text-vision scores to retain visual tokens, those near the sequence end often get higher weights. In 2D image space, this biases retention toward the lower half or bottom-right corner. Clearly, it’s unrealistic to assume the lower half of all images is more important. Such bias can significantly hurt compression performance. As shown in Figure 2, a comprehensive review and comparison of key studies jiang2025kind ; wen2025stop ; wen2025token ; Zhang2024:SparseVLM ; liu2024multi ; liu2025vidcom2 confirms our hypothesis: even well-crafted attention-based methods can underperform simple random pruning or pooling. Detailed analysis is in Appendix B.
Inherent Limitations of Token Compression. Beyond performance degradation from methodological design, does token compression face inherent limitations? Is it universally applicable across various tasks? For multi-modal large language models, wen2025token shows most existing compression methods underperform on visual grounding tasks, with significant drops on benchmarks like RefCOCO yu2016modeling . In OCR-related parsing yang2024cc ; ouyang2024omnidocbench , documents with dense layouts yield highly information-rich visual tokens. Compressing these risks severe information loss and degraded performance. Beyond vision, current methods also face inherent limits in other modalities. In Automatic Speech Recognition (ASR) and Automatic Speech Translation (AST) ardila2019common ; conneau2023fleurs , audio is encoded and then decoded into text using an MLLM abouelenin2025phi ; chu2024qwen2 . Audio tokens in ASR and AST are dense and temporally continuous. Pruning or merging them disrupts this continuity, leading to fragmented recognition or translation. Similarly, translation across languages in the text modality may suffer significant degradation under high compression ratios.
@@ 4.3 Fair Comparison
Rethinking FLOPs and Compression Ratios as Efficiency Metrics. Many token compression methods report speedup by estimating FLOPs reductions or directly using token compression ratios. But do FLOPs or compression ratios truly reflect real acceleration? Our analysis shows that, even with similar compression ratios or FLOPs, methods often vary significantly in runtime latency. Investigating further, we find: (i) Importance-based compression often uses attention scores Chen2024:FastV ; zhang2024sparsevlm , but this can limit compatibility with efficient attention mechanisms like Flash Attention dao2022flashattention ; dao2023flashattention2 , potentially contributing to the substantial discrepancy between FLOPs and actual runtime latency observed in some approaches. (ii) Some methods pursue high compression via progressive compression across layers xing2024pyramiddrop , adding overhead that offsets the gains from token reduction. Thus, we argue that runtime latency should be prioritized in evaluations, as FLOP or token count reductions don’t always yield real-world speedup.
@@
Token Compression Evaluation: The Benchmarking Gap. Current token compression methods are primarily evaluated using general-purpose benchmarks, which are not tailored to capture the unique challenges of token compression. As a result, some benchmarks—such as ScienceQA lu2022learn and VizWiz bigham2010vizwiz —paradoxically show improved performance under certain compression settings, or minimal degradation across varying compression ratios. These observations defy intuition and suggest that existing benchmarks may fail to meaningfully reflect the trade-offs introduced by token compression. This discrepancy raises concerns about the validity of current evaluation practices. In particular, benchmarks that fail to penalize information loss introduced by compression may obscure meaningful differences between methods. Additionally, the lack of task diversity and compression-sensitive metrics further limits our understanding of how these methods behave in realistic scenarios. Without dedicated benchmarks, it remains unclear whether observed gains are due to genuine improvements in compression quality or artifacts of misaligned evaluation settings.
@@ 5.2 Dedicated Benchmarks for Token Compression
Given the current limitations in evaluating data-centric token compression methods using general-purpose benchmarks, we envision the development of a dedicated benchmark specifically designed to evaluate them. Such a benchmark should comprehensively span diverse domains—including natural language processing, computer vision, and multi-modal tasks—and incorporate task-specific challenges particularly relevant to token compression, such as optical character recognition (OCR) parsing ouyang2024omnidocbench ; zhang2024ocr and automatic speech recognition (ASR) conneau2023fleurs ; park2024long . Furthermore, it is essential that this benchmark jointly considers both task performance and latency, as both are critical for real-world deployment. A well-rounded benchmark of this nature would enable more rigorous, fair, and holistic evaluation of token compression techniques, ultimately driving progress in this area.
@@ 6.1 Model-Centric Compression as a Superior Alternative
Model-centric compression methods, such as pruning Network_Pruning ; JMLR:v12:huang11b ; Deep_Compression ; Rao2021:DynamicViT ; jang2017Gumbel-Softmax ; Pan2021:IA-RED2 , quantization rokh2023comprehensive ; Wang:HAQ ; zhou2018adaptive ; yang2019quantization ; Lin:AWQ ; Frantar:GPTQ , and knowledge distillation Hinton:KD ; zhang2019self_distillation ; zhang2021self_distillation ; park2019relational_distillation , have long been established as effective techniques for reducing model size and computational cost. Proponents argue that this paradigm is reliable for deployment in resource-constrained environments and maintains performance consistency. For example, pruning techniques such as DynamicViT Rao2021:DynamicViT dynamically remove uninformative tokens during inference, reducing the computational load by up to 30–40% with minimal impact on accuracy. Proponents of this view claim that this approach achieves substantial speedups without discarding any original data. In contrast, data-centric methods that prune input tokens risk removing critical contextual information, which may degrade performance.
Counterargument. Although model-centric compression is effective, it faces scalability issues as models and datasets grow, requiring costly full retraining and processing of entire inputs. In contrast, data-centric compression reduces input complexity upfront, easing computational burdens. Some data-centric methods update only a small parameter subset Zhao2024:DyT ; liu2024Sparse-Tuning ; Lee2024:DETM , while others enable training-free deployment bolya2023tome ; Chen2024:FastV ; li2024snapkv . Combining both approaches can improve efficiency without sacrificing accuracy azeemi2023data ; kousar2025pruning , making data-centric methods a complement to model-centric techniques.
@@ 6.2 Advanced Model Architectures as a more Promising Direction
Another argument against data-centric compression is the continued advancement of model architectures that can inherently handle large datasets and long sequences more efficiently gu2021efficiently ; goldstein2024goldfinch ; peng2023rwkv . The development of transformer-based architectures, such as Vision Transformers ViT , Swin Transformers liu2021swin , and large language models like GPT-3 brown2020language , has shown significant improvements in both accuracy and scalability. These architectures integrate advanced techniques, such as hierarchical processing, self-attention mechanisms, and dynamic sparsity, enabling them to process large amounts of data efficiently. For example, Swin Transformers liu2021swin utilize a window-based self-attention mechanism, which reduces the computational complexity of the standard attention mechanism, making it feasible to scale models to much larger datasets and sequences. Proponents of this view argue that as these advanced models continue to evolve, there may be less need for aggressive input compression, as these models are inherently better equipped to handle large-scale data directly.
Counterargument. Advanced model architectures offer strong performance but demand substantial computational resources, especially during training llama4 ; yang2025Qwen3 ; OpenAI:GPT-4 . Data-centric compression reduces computational load early by simplifying input data, enabling more efficient training and inference without sacrificing accuracy. Techniques like token pruning and augmentation preserve or improve performance by focusing on informative data. Combined with advanced architectures, data-centric methods enhance efficiency and maintain high performance gao2020pile ; wettig2024qurating , making them complementary rather than competitive.